ZETA FUNCTIONS OF FINITE GRAPHS AND COVERINGS

H. M. Stark A. A. Terras

Abstract. Three different zeta functions are attached to finite, connected, possibly irregular graphs. They originate with a zeta function of Ihara which is an analogue of Riemann’s as well as Selberg’s zeta function. The three zeta functions are associated to one vertex variable, two variables for each edge, and 2r(2r − 1) path variables, respectively. Here r is the number of generators of the fundamental group of X. We show how to specialize the variables of the last two zeta functions to obtain the first and we give elementary proofs of generalizations of Ihara’s formula saying that the zeta function for a regular graph is the reciprocal of a polynomial. Many examples of covering graphs are considered.

Typeset by AMS-TEX